OpenStudy (anonymous):

what is \[\int\limits_{-\infty}^{0} dx/(8-3x)^{5/3}\] ?

OpenStudy (anonymous):

please help..

OpenStudy (anonymous):

rewrite it as \[\lim_{t\rightarrow_-\infty}{\int_t^0{\frac{dx}{(8-3x)^{\frac{5}{3}}}}}\]

OpenStudy (anonymous):

use substitution :-D

OpenStudy (anonymous):

ok.=) sorry. thanks..

OpenStudy (anonymous):

so i'll use 8-3x as u?

OpenStudy (anonymous):

1/36

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so when i solve, what is the value of the negative infinity?

OpenStudy (anonymous):

Just find the limit of the expression, as t goes to - infinity

OpenStudy (anonymous):

let 8-3x = u du/dx = -3 du/-3 = dx the integral becomes :: -1/3 integral (u^(-5/3)du) the answer to integral (u^(-5/3)du) -> u^(-2/3) / (-2/3) -1/3 * ( -3u^(-2/3)/2) u^(-2/3)/2 1/(2* (8-3x)^(2/3)) at 0 its : 1/(2*4) = 1/8 at -infinity its : 0.5lim(1/(8-3x)^(2/3)) = 0 so its 1/8 i hope didnt make any mistakes

OpenStudy (anonymous):

why is it at 0.5?

OpenStudy (anonymous):

i took the 1/2 out of the limit

OpenStudy (anonymous):

Coolsector is right. he just took out the 0.5 factor from the expression :-D

OpenStudy (anonymous):

ah. i see.=) thank you guys.